# Solitons in binary compounds with stacked two-dimensional honeycomb lattices

### v1

@@ Abstract
We model the electronic properties of thin films of binary compounds with stacked layers where each layer is a two-dimensional honeycomb lattice with two atoms per unit cell. The two atoms per cell are assigned different onsite energies in order to consider six different stacking orders: ABC, ABA, AA, ${}^{\prime}$, ${}^{\prime}$, and ${}^{\prime}$. Using a minimal tight-binding model with nearest-neighbor hopping, we consider whether a fault in the texture of onsite energies in the vertical, stacking direction supports localized states, and we find localized states within the bulk band gap for ABC, ABA, and ${}^{\prime}$ stacking. Depending on the stacking type, parameter values, and whether the soliton is atomically sharp or a smooth texture, there are a range of different band structures including soliton bands that are either isolated or that hybridize with other states, such as surface states, and soliton bands that are either dispersive or flat, the latter yielding narrow features in the density of states. We discuss the relevance of our results to specific materials including graphene, hexagonal boron nitride and other binary compounds.
@@ I Introduction
Figure 1: Summary of the main results for the six different stacking types. The first row (a)-(f) shows the faultless systems with a schematic side view of the stacking for $N=4$ layers with red (yellow) circles indicating A (B) atoms, horizontal (vertical) solid lines indicating nearest-neighbor intralayer (interlayer) hopping. At the bottom of each panel is the band structure for faultless stacking for $N=8$ layers near low energy as a function of wave vector component $k_{x}$ (with $k_{y}=0$) measured from the center of the $K_{+}$ valley, where $k_{c}=\gamma_{1}/(\hbar v)$. The second row (g)-(l) shows the lattices for $N=4$ layers with an atomically-sharp fault in the onsite energies occurring between the middle two layers, denoted $(2,2)$, and the bottom of each panel shows the band structures for $N=8$ layers with a fault in the onsite energies occurring between the middle two layers, denoted $(4,4)$. Horizontal arrows show the energies at $k_{x}=0$ of the states localized on the soliton. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [35], $a=2.46$Å [36], and the magnitude of the alternating onsite energies is $u=U/\gamma_{1}=0.4$.
@@
[…]
@@
We focus on electronic properties near the corner of the first Brillouin zone (K point) and Fig. 1 summarizes our results. The top row shows the six faultless systems with their electronic band structures. Three of them (ABC, ABA, and ${}^{\prime}$ stacking) exhibit band gaps, and these are the configurations which support localized states on a stacking fault that lie within the band gap. A sharp stacking fault consists of an inversion of the signs of the onsite energy $U$ between two layers without any change in the intra- or interlayer hopping, as shown in the bottom row of Fig. 1; we refer to this as a soliton because it creates a change in texture of the onsite energies either side of the fault. We find a range of different behaviors, depending on the stacking; arrows in Fig. 1 (bottom row) show the energies of states localized on the soliton exactly at the K point. For ABC stacking, there are two localized states. One of them is within the valence band, but another generally lies within the bulk band gap. Depending on parameter values, it may hybridize with surface states, as shown in Fig. 1(g). For ${}^{\prime}$, there are also two localized states, but they don’t lie near zero energy; the two states within the band gap in Fig. 1(h) are surface states. For ABA stacking, there are two localized states. One of them is within the conduction band, but another generally lies within the bulk band gap, Fig. 1(i), and there are no surface states to hybridize with. As this band eventually moves into the conduction band, we refer to the corresponding texture as an antisoliton whereas the band shown in Fig. 1(g) for ABC stacking moves into the valence band, and its texture is called a soliton. In general, both of these stackings (ABC and ABA) can support either a soliton or an antisoliton, depending on the position of the fault.
@@
For both ${}^{\prime}$ and AA stacking, there are no states localized on the fault, although the presence of the fault does have an impact on the band structure, Fig. 1(j) and Fig. 1(k), respectively. For ${}^{\prime}$, a single fault supports four localized states: Two of them are in the bulk bands, but two of them lie within the bulk band gap and, depending on parameter values, they may hybridize with each other as shown in Fig. 1(l).
@@
In the following, we describe the cases shown in Fig. 1, focusing on the systems with a single, isolated soliton band, and modeling their parameter dependence. We describe the methodology used in Section II. We describe rhombohedral stacking (ABC and ${}^{\prime}$) in Section III, Bernal stacking (ABA and ${}^{\prime}$) in Section IV, and AA stacking (AA and ${}^{\prime}$) in Section V. The properties of smooth solitons with a finite width are described in Section VI, and defects which consist of a reversal of the signs of the onsite energies on a single layer only [33] are described in Section VII. Section VIII briefly describes the robustness of soliton features in the density of states to the presence of interlayer disorder. Relevance to particular materials and the stability of structural defects is discussed in Section IX.
@@ II Methodology
For the band structure plots we numerically diagonalize the Hamiltonian (6). We use tight-binding parameters experimentally measured [35] for Bernal-stacked bilayer graphene $\gamma_{0}=3.16\,$eV and $\gamma_{1}=0.381\,$eV, and lattice constant $a=2.46$Å [36]. Band structures are calculated in the vicinity of valley $K_{+}$ with wave vector ${\mathbf{K_{+}}}=(4\pi/(3a),0)$ by shifting the wave vector as ${\mathbf{q}}={\mathbf{K_{+}}}+{\mathbf{k}}$, with ${\mathbf{k}}=(k_{x},k_{y})$. Specifically, we determine the energy eigenvalues $E_{n}$ on a square grid of points centered on $K_{+}$. The band structures have valley degeneracy (under a suitable rotation) and they are approximately isotropic around each valley at the energies we consider, so we plot them as a function of component $k_{x}$ (with $k_{y}=0$), normalized by the characteristic wave vector $k_{c}=\gamma_{1}/\hbar v$ [37] where $v=\sqrt{3}a\gamma_{0}/(2\hbar)$ is the velocity related to intralayer hopping. The density of states $g(E)$ per unit energy per unit area ($L^{2}$) is determined numerically by approximation using a Lorentzian with a finite width $\delta$,
@@ III.1.2 Single soliton
For finite $U$, we consider a sharp fault consisting of two adjacent sites with negative onsite energies as shown in Fig. 1(g). Band structures and density of states for $N=16$ layers and a sharp fault occurring between the middle two layers, denoted $(8,8)$, are shown in Fig. 2. There are two bands localized on the fault, as indicated by the arrows in the band structure plots of Fig. 2. Both of these bands move into the valence band for $k_{x}\gg k_{c}$, hence we refer to this fault as a soliton [48]. We focus on the higher energy band of the two because it is near zero energy. For $U\leq\gamma_{1}/2$, this soliton band hybridizes with the flat bands arising from the surface states, Fig. 2(b) and Fig. 2(c). Note that the soliton changes the texture of onsite energies, flipping the sign of the onsite energy on the top surface (site $B_{N}$) so that both surface bands move into the conduction band for $k_{x}\gg k_{c}$.
@@
Note that expressions (53,54) for the energy $E_{\mathrm{sol}}$ of the band localized on the soliton are independent of the position of the soliton and the number of layers $N$ in the system.
At $k=0$, the soliton is localized on the two dimer sites $B_{m}$ and $A_{m+1}$, i.e. it is given by $\psi_{\mathrm{sol}}=(\psi_{B_{m}}+\psi_{A_{m+1}})/\sqrt{2}$ where $\psi_{j}$ denotes the atomic orbital on site $j$. For $U\approx\gamma_{1}/2$ and $k\ll k_{c}$ we derive an effective $3\times 3$ Hamiltonian describing the hybridization of the soliton state, $\psi_{\mathrm{sol}}=(\psi_{B_{m}}+\psi_{A_{m+1}})/\sqrt{2}$, with the two surface states $\psi_{A_{1}}$ and $\psi_{B_{N}}$. This is done using the linear expansion $f({\mathbf{K_{\xi}}}+{\bf k})\approx-\sqrt{3}a(\xi k_{x}-ik_{y})/2$, eliminating the other orbitals [49, 34] and performing an expansion for $k/k_{c}\ll 1$ and $|U/\gamma_{1}-1/2|\ll 1$. In basis $\psi_{A_{1}},(\psi_{B_{m}}+\psi_{A_{m+1}})/\sqrt{2},\psi_{B_{N}}$, the effective Hamiltonian is
@@ III.2.1 Faultless system
Figure 6: Low-energy band structures for ${}^{\prime}$ stacking. Each column shows a different value of the onsite energies $U$, each row shows a different system. The top row shows a faultless system with $N=16$ layers, whereas the second row shows $N=16$ layers with a sharp fault at its center $(8,8)$. The third row shows a faultless system with $N=15$ layers, and the final row shows $N=15$ layers with a sharp fault near its center $(8,7)$. Horizontal arrows show the energies at $k_{x}=0$ of the states localized on the soliton. Note that the axis scales for the final column (with $U=2.00\gamma_{1}$) differ from the rest. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [35], $a=2.46$Å [36].
For bulk, faultless ${}^{\prime}$ stacking, the unit cell consists of four atoms such as $A_{1}$, $B_{1}$, $A_{2}$, $B_{2}$ (in Fig. 1(b)), and the Hamiltonian may be written as
@@ III.2.2 Single soliton
The band structure for ${}^{\prime}$ stacking depends on whether there is an even or odd number of layers, even for a faultless system. The band structure for a faultless system with an odd number of layers $N$ is shown in Fig. 6 (third row). For $N\geq 3$, at $k=0$, the $N$ bands are at six distinct energies with $E=+U$ and $E=-U$ levels arising from the surface states, and from the dimers: $E=+U+\gamma_{1}$, $E=+U-\gamma_{1}$, $E=-U+\gamma_{1}$, and $E=-U-\gamma_{1}$, each of these occurring $(N-1)/2$ times. As with the faultless even $N$ case, the surface bands appear at lower energy ($E=\pm U$) than the bulk bands for $20\leq U<\gamma_{1}/$ 2, and the bulk bands touch at $k=0$ and $U=\gamma_{1}$. Nevertheless, the band structure is more similar to that of the even $N$ case with a fault. This can be seen by considering the bands for nonzero $k$ values: For $0<U<\gamma_{1}$, there is a small anticrossing between two low-energy bands, then, for $U=\gamma_{1}$, the bands touch, and, for $U>\gamma_{1}$, there is a band gap $E_{{\mathrm{g}}}=2(U-\gamma_{1})$.
@@ IV.1.1 Faultless system
For a thin film of finite thickness, a faultless system is gapped with bandgap $E_{\mathrm{g}}\approx 2U$, Fig. 1(c). This can be understood by considering $k=0$ where the intralayer hopping is zero and the system separates into isolated atoms with energies $E=\pm U$ plus a $N$-mer which is a single chain of $N$ atoms ($N=4$ in Fig. 1(c)). This chain corresponds to the CDW model [41, 42, 15, 43, 44, 16] which is a one-dimensional chain with constant hopping $\gamma_{1}$ and alternating onsite energies ±$\pm U$, having a band gap of $2U$, as described in the methodology, Section II.
@@ IV.1.2 Single soliton
For a system with a single fault, there are two bands localized on the fault, one of which generally lies within the bulk band gap, Fig. 1(i) and Fig. 7. The nature of this band depends on whether the fault occurs after an even or odd number of layers [48] as shown in Fig. 7. When the fault occurs after an even number of layers, Fig. 1(i), Fig. 7(a), and Fig. 7(b), the fault consists of two consecutive onsite energies of $+U$ (on atoms $A_{2}$ and $B_{3}$ in Fig. 1(i)) and we refer to this as an antisoliton; the energy band moves into the conduction band for $k\gg k_{c}$. However, when the fault occurs after an odd number of layers, Fig. 7(c) and Fig. 7(d), the fault consists of two consecutive onsite energies of $-U$, a soliton texture, and the energy band moves into the valence band for $k\gg k_{c}$.
We consider $k=0$ where the intralayer hopping is zero. Once again, the system separates into isolated atoms with energies $E=\pm U$ plus a $N$-mer ($N=4$ in Fig. 1(i)). But now the $N$-mer, which corresponds to the CDW model [41, 42, 15, 43, 44, 16], has a soliton or antisoliton in it [42, 44]. It was shown [44] that the energy of the state localized on such a fault lies within the bulk band gap as long as $0<\gamma_{1}/U<N$, when the fault lies at the center of the system of $N$ layers.
@@
For $k=0$, the soliton state is localized on the soliton [42, 44], and, for $k_{x}=k_{c}$ (and $k_{y}=0$), the state is still localized near the soliton as shown in Fig. 4(c). Since the system is equivalent to the CDW model at $k=0$, the soliton state is polarized, with maximum polarization at $k=0$, Fig. 4(d). In this case, the equivalence with the CDW model only applies to the $N$-mer (the atomic sites connected by interlayer hopping $\gamma_{1}$), excluding the isolated non-$N$-mer atoms. Therefore, we modify the definition of the translation operator $T_{a/2}$, Eq. (47), in order to describe translation between the atoms on the $N$-mer only,

### v2

@@ Abstract
We model the electronic properties of thin films of binary compounds with stacked layers where each layer is a two-dimensional honeycomb lattice with two atoms per unit cell. The two atoms per cell are assigned different onsite energies in order to consider six different stacking orders: ABC, ABA, AA, ABC, ABA, and AA. Using a minimal tight-binding model with nearest-neighbor hopping, we consider whether a fault in the texture of onsite energies in the vertical, stacking direction supports localized states, and we find localized states within the bulk band gap for ABC, ABA, and AA stacking. Depending on the stacking type, parameter values, and whether the soliton is atomically sharp or a smooth texture, there are a range of different band structures including soliton bands that are either isolated or that hybridize with other states, such as surface states, and soliton bands that are either dispersive or flat, the latter yielding narrow features in the density of states. We discuss the relevance of our results to specific materials including graphene, hexagonal boron nitride and other binary compounds.
@@ I Introduction
Figure 1: Summary of the main results for the six different stacking types. The first row (a)-(f) shows the faultless systems with a schematic side view of the stacking for $N=8$ layers with red (yellow) circles indicating A (B) atoms, horizontal (vertical) solid lines indicating nearest-neighbor intralayer (interlayer) hopping. The bottom of each panel shows the corresponding band structure near low energy as a function of wave vector component $k_{x}$ (with $k_{y}=0$) measured from the center of the $K_{+}$ valley, where $k_{c}=\gamma_{1}/(\hbar v)$. The second row (g)-(l) shows the lattices for $N=8$ layers with an atomically-sharp fault in the onsite energies occurring between the middle two layers and the bottom of each panel shows the corresponding band structures. Horizontal arrows show the energies at $k_{x}=0$ of the states localized on the soliton. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [40], $a=2.46$Å [41], and the magnitude of the alternating onsite energies is $u=U/\gamma_{1}=0.4$.
@@
In this paper, we investigate the possible existence of a single, isolated band localized on a stacking fault in quasi-2D materials consisting of layers of atoms on a honeycomb lattice. We consider materials with two different chemical elements such as hexagonal boron nitride [26, 27], using a minimal tight-binding model to provide a generic description, modeling stacking faults for different stacking orders. Specifically, for each layer, we consider two nonequivalent atoms per unit cell, labeled A and B, each with a single orbital which is isotropic within the plane, e.g. a $p_{z}$ orbital for sp2 hybridization in graphene or hexagonal boron nitride. The two atoms have different onsite energies $\epsilon_{A}=U$ and $\epsilon_{B}=-U$ such that $U=(\epsilon_{A}-\epsilon_{B})/2$ and $\epsilon_{A}+\epsilon_{B}=0$. We consider naturally occurring stackings, i.e., those whereby every atom is either directly above or below another atom, or above or below the center of a hexagon. There are six of them [28, 30, 31, 29], as shown in the top row of Fig. 1, consisting of rhombohedral stacking (ABC), Bernal stacking (ABA), and AA stacking. For each of these, we also include a primed version (ABC, ABA, and AA) whereby every other layer has the sign of the onsite energy $U$ reversed [28, 29].
For rhombohedral stacking (ABC) with alternating onsite energies, dimensional reduction [17], by considering the in-plane wave vector as a fixed parameter, relates the Hamiltonian to that of the Rice-Mele model [32], which is a generalization of the SSH model with alternating onsite energies as well as alternating hopping in one dimension. This is generally not topological, but, at certain wave vectors ($k_{x}=k_{c}$ in Fig. 1), the intra- and interlayer hoppings are effectively equal, and ABC stacking is related to the charge density wave (CDW) model [33, 34, 15, 35, 36, 16] which is a one-dimensional $\mathbb{Z}_{2}$ topological insulator [37] with constant hoppings and alternating onsite energies. Stacking faults for ABC stacking are analogous to solitons in the CDW model [33, 34, 36], albeit only at a certain wave vector, and our aim is to explore this analogy.
@@
We focus on electronic properties near the corner of the first Brillouin zone (K point) and Fig. 1 summarizes our results. The top row shows the six faultless systems with their electronic band structures. Three of them (ABC, ABA, and AA stacking) exhibit band gaps, and these are the configurations which support localized states on a stacking fault that lie within the band gap. A sharp stacking fault consists of an inversion of the signs of the onsite energy $U$ between two layers without any change in the intra- or interlayer hopping, as shown in the bottom row of Fig. 1; we refer to this as a soliton because it creates a change in texture of the onsite energies either side of the fault. We find a range of different behaviors, depending on the stacking; arrows in Fig. 1 (bottom row) show the energies of states localized on the soliton exactly at the K point. For ABC stacking, there are two localized states. One of them is within the valence band, but another generally lies within the bulk band gap. Depending on parameter values, it may hybridize with surface states, as shown in Fig. 1(g). For ABC, there are also two localized states, but they don’t lie near zero energy; the two states within the band gap in Fig. 1(h) are surface states. For ABA stacking, there are two localized states. One of them is within the conduction band, but another generally lies within the bulk band gap, Fig. 1(i), and there are no surface states to hybridize with. As this band eventually moves into the conduction band, we refer to the corresponding texture as an antisoliton whereas the band shown in Fig. 1(g) for ABC stacking moves into the valence band, and its texture is called a soliton. In general, both of these stackings (ABC and ABA) can support either a soliton or an antisoliton, depending on the position of the fault.
@@
For both ABA and AA stacking, there are no states localized on the fault, although the presence of the fault does have an impact on the band structure, Fig. 1(j) and Fig. 1(k), respectively. For AA, a single fault supports four localized states: Two of them are in the bulk bands, but two of them lie within the bulk band gap and, depending on parameter values, they may hybridize with each other as shown in Fig. 1(l).
@@
In the following, we describe the cases shown in Fig. 1, focusing on the systems with a single, isolated soliton band, and modeling their parameter dependence. We describe the methodology used in Section II. We describe rhombohedral stacking (ABC and ABC) in Section III, Bernal stacking (ABA and ABA) in Section IV, and AA stacking (AA and AA) in Section V. The properties of smooth solitons with a finite width are described in Section VI, and defects which consist of a reversal of the signs of the onsite energies on a single layer only [38] are described in Section VII. Section VIII briefly describes the robustness of soliton features in the density of states to the presence of interlayer disorder. In Section IX, we describe ballistic, coherent electronic transport for ABC stacking, comparing the energy dependence of conductivity for systems with and without a soliton. Section X describes the influence of additional tight-binding parameters beyond the minimal model, which only has nearest-neighbor intra- and interlayer coupling. Relevance to particular materials and the stability of structural defects is discussed in Section XI.
@@ II Methodology
For the band structure plots we numerically diagonalize the Hamiltonian (1). We use tight-binding parameters experimentally measured [40] for Bernal-stacked bilayer graphene $\gamma_{0}=3.16\,$eV and $\gamma_{1}=0.381\,$eV, and lattice constant $a=2.46$Å [41]. Band structures are calculated in the vicinity of valley $K_{+}$ with wave vector ${\mathbf{K_{+}}}=(4\pi/(3a),0)$ by shifting the wave vector as ${\mathbf{q}}={\mathbf{K_{+}}}+{\mathbf{k}}$, with ${\mathbf{k}}=(k_{x},k_{y})$. Specifically, we determine the energy eigenvalues $E_{n}$ on a square grid of points centered on $K_{+}$. The band structures have valley degeneracy (under a suitable rotation) and they are approximately isotropic around each valley at the energies we consider, so we plot them as a function of component $k_{x}$ (with $k_{y}=0$), normalized by the characteristic wave vector $k_{c}=\gamma_{1}/\hbar v$ [42] where $v=\sqrt{3}a\gamma_{0}/(2\hbar)$ is the velocity related to intralayer hopping. By plotting $k_{x}$ as a function of $k_{c}$ and energy $E$ as a function of $\gamma_{1}$, there are only two free parameters in our model, namely the ratios $\gamma_{0}/\gamma_{1}$ and $U/\gamma_{1}$. The density of states $g(E)$ per unit energy per unit area ($L^{2}$) is determined numerically by approximation using a Lorentzian with a finite width $\delta$,
@@
States localized on solitons may or may not appear in the bulk band gap, but we focus on those that do. Such states are extended states in the in-plane direction, but are localized on the soliton in the out-of-plane ($z$) direction. We identify them by observing the form of the state in the atomic basis of the Hamiltonian (1); an example is shown in Fig. 4(a).
@@ III.1.2 Single soliton
For finite $U$, we consider a sharp fault consisting of two adjacent sites with negative onsite energies as shown in Fig. 1(g). Band structures and density of states for $N=16$ layers and a sharp fault occurring between the middle two layers, denoted $(8,8)$, are shown in Fig. 2. There are two bands localized on the fault, as indicated by the arrows in the band structure plots of Fig. 2. Both of these bands move into the valence band for $k_{x}\gg k_{c}$, hence we refer to this fault as a soliton [47]. We focus on the higher energy band of the two because it is near zero energy. For $U\leq\gamma_{1}/2$, this soliton band hybridizes with the flat bands arising from the surface states, Fig. 2(b) and Fig. 2(c). Note that the soliton changes the texture of onsite energies, flipping the sign of the onsite energy on the top surface (site $B_{N}$). This can be seen in Fig. 1(g) where the soliton (between sites $B_{4}$ and $A_{5}$) changes the order of the onsite energies for higher layers (with layer index $j\geq 5$) so that the final orbital, $B_{8}$, has the same onsite energy as the first, $A_{1}$, in contrast to the faultless case, Fig. 1(a). As a result, both surface states, localized on $A_{1}$ and $B_{8}$, are at positive energy and move into the conduction band for $k_{x}\gg k_{c}$.
@@
Note that expressions (17,18) for the energy $E_{\mathrm{sol}}$ of the band localized on the soliton are independent of the position of the soliton and the number of layers $N$ in the system.
Soliton states are extended states in the in-plane direction, but are localized in the out-of-plane ($z$) direction. At $k=0$, the soliton state is localized on the two dimer sites $B_{m}$ and $A_{m+1}$, i.e. it is given by $\psi_{\mathrm{sol}}=(\psi_{B_{m}}+\psi_{A_{m+1}})/\sqrt{2}$ where $\psi_{j}$ denotes the atomic orbital on site $j$. For $U\approx\gamma_{1}/2$ and $k\ll k_{c}$ we derive an effective $3\times 3$ Hamiltonian describing the hybridization of the soliton state, $\psi_{\mathrm{sol}}=(\psi_{B_{m}}+\psi_{A_{m+1}})/\sqrt{2}$, with the two surface states $\psi_{A_{1}}$ and $\psi_{B_{N}}$. This is done using the linear expansion $f({\mathbf{K_{\xi}}}+{\bf k})\approx-\sqrt{3}a(\xi k_{x}-ik_{y})/2$, eliminating the other orbitals [48, 39] and performing an expansion for $k/k_{c}\ll 1$ and $|U/\gamma_{1}-1/2|\ll 1$. In basis $\psi_{A_{1}},(\psi_{B_{m}}+\psi_{A_{m+1}})/\sqrt{2},\psi_{B_{N}}$, the effective Hamiltonian is
@@ III.2.1 Faultless system
For bulk, faultless ABC stacking, the unit cell consists of four atoms such as $A_{1}$, $B_{1}$, $A_{2}$, $B_{2}$ (in Fig. 1(b)), and the Hamiltonian may be written as
@@
Figure 6: Low-energy band structures for ABC stacking. Each column shows a different value of the onsite energies $U$, each row shows a different system. The top row shows a faultless system with $N=16$ layers, whereas the second row shows $N=16$ layers with a sharp fault at its center $(8,8)$. The third row shows a faultless system with $N=15$ layers, and the final row shows $N=15$ layers with a sharp fault near its center $(8,7)$. Horizontal arrows show the energies at $k_{x}=0$ of the states localized on the soliton. Note that the axis scales for the final column (with $U=2.00\gamma_{1}$) differ from the rest. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [40], $a=2.46$Å [41].
@@ III.2.2 Single soliton
The band structure for ABC stacking depends on whether there is an even or odd number of layers, even for a faultless system. The band structure for a faultless system with an odd number of layers $N$ is shown in Fig. 6 (third row). For $N\geq 3$, at $k=0$, the $N$ bands are at six distinct energies with $E=+U$ and $E=-U$ levels arising from the surface states, and from the dimers: $E=+U+\gamma_{1}$, $E=+U-\gamma_{1}$, $E=-U+\gamma_{1}$, and $E=-U-\gamma_{1}$, each of these occurring $(N-1)/2$ times. As with the faultless even $N$ case, the surface bands appear at lower energy ($E=\pm U$) than the bulk bands for $20\leq U<\gamma_{1}/$ 2, and the bulk bands touch at $k=0$ and $U=\gamma_{1}$. Nevertheless, the band structure is more similar to that of the even $N$ case with a fault. This can be seen by considering the bands for nonzero $k$ values: For $0<U<\gamma_{1}$, there is a small anticrossing between two low-energy bands, then, for $U=\gamma_{1}$, the bands touch, and, for $U>\gamma_{1}$, there is a band gap $E_{{\mathrm{g}}}=2(U-\gamma_{1})$.
@@ IV.1.1 Faultless system
For a thin film of finite thickness, a faultless system is gapped with bandgap $E_{\mathrm{g}}\approx 2U$, Fig. 1(c). This can be understood by considering $k=0$ where the intralayer hopping is zero and the system separates into isolated atoms with energies $E=\pm U$ plus a $N$-mer which is a single chain of $N$ atoms ($N=8$ in Fig. 1(c)). This chain corresponds to the CDW model [33, 34, 15, 35, 36, 16] which is a one-dimensional chain with constant hopping $\gamma_{1}$ and alternating onsite energies ±$\pm U$, having a band gap of $2U$, as described in the methodology, Section II.
@@ IV.1.2 Single soliton
For a system with a single fault, there are two bands localized on the fault, one of which generally lies within the bulk band gap, Fig. 1(i) and Fig. 7. The nature of this band depends on whether the fault occurs after an even or odd number of layers [47] as shown in Fig. 7. When the fault occurs after an even number of layers, Fig. 1(i), Fig. 7(a), and Fig. 7(b), the fault consists of two consecutive onsite energies of $+U$ (on atoms $A_{4}$ and $B_{5}$ in Fig. 1(i)) and we refer to this as an antisoliton; the energy band moves into the conduction band for $k\gg k_{c}$. However, when the fault occurs after an odd number of layers, Fig. 7(c) and Fig. 7(d), the fault consists of two consecutive onsite energies of $-U$, a soliton texture, and the energy band moves into the valence band for $k\gg k_{c}$.
We consider $k=0$ where the intralayer hopping is zero. Once again, the system separates into isolated atoms with energies $E=\pm U$ plus a $N$-mer ($N=8$ in Fig. 1(i)). But now the $N$-mer, which corresponds to the CDW model [33, 34, 15, 35, 36, 16], has a soliton or antisoliton in it [34, 36]. It was shown [36] that the energy of the state localized on such a fault lies within the bulk band gap as long as $0<\gamma_{1}/U<N$, when the fault lies at the center of the system of $N$ layers.
@@
Soliton states are extended states in the in-plane direction, but are localized in the out-of-plane ($z$) direction. For $k=0$, the soliton state is localized on the soliton [34, 36], and, for $k_{x}=k_{c}$ (and $k_{y}=0$), the state is still localized near the soliton as shown in Fig. 4(c). Since the system is equivalent to the CDW model at $k=0$, the soliton state is polarized, with maximum polarization at $k=0$, Fig. 4(d). In this case, the equivalence with the CDW model only applies to the $N$-mer (the atomic sites connected by interlayer hopping $\gamma_{1}$), excluding the isolated non-$N$-mer atoms. Therefore, we modify the definition of the translation operator $T_{a/2}$, Eq. (13), in order to describe translation between the atoms on the $N$-mer only,